In a scalene triangle, if the shortest side has a measurement of 10 units long,what are the lengths of the other 2 sides. Information given:
Short side s-3.5=10 Base s+7.5 opposite side s+6

Answer by ewatrrr(24785)

 
Hi,
Short side s-3.5=10 ⇒ s = 13.5
Base s+7.5 , 13.5 + 7.5 = 21 units
opposite side s+6 , 13.5 + 6 = 19.5 units